# Multi-Preconditioned LBFGS for Training Finite-Basis PINNs

### v1

@@
11institutetext: Marc Salvadó-Benasco 22institutetext: Università della Svizzera Italiana; e-mail: 22email: marc.salvado@usi.ch 33institutetext: Aymane Kssim 44institutetext: Toulouse INP-ENSEEIHT, IRIT, ANITI; 44email: aymane.kssim@toulouse-inp.fr 55institutetext: Alexander Heinlein 66institutetext: Delft Institute of Applied Mathematics, TU Delft; 66email: a.heinlein@tudelft.nl 77institutetext: Rolf Krause 88institutetext: King Abdullah University of Science and Technology; 88email: rolf.krause@kaust.edu.sa 99institutetext: Serge Gratton 1010institutetext: Toulouse INP-ENSEEIHT, IRIT, ANITI; 1010email: serge.gratton@toulouse-inp.fr 1111institutetext: Alena Kopaničáková 1212institutetext: Toulouse INP-ENSEEIHT, IRIT, ANITI; 1212email: alena.kopanicakova@toulouse-inp.fr
@@ 1 Introduction
This manuscript is organized as follows. Section 2 introduces the FBPINN architecture, Section 3 presents the multi-preconditioned LBFGS algorithm, and numerical results and conclusions are given in Section 4 and Section 5, respectively.
@@ 2 Finite Basis PINNs (FBPINNs)
Figure 1: Window functions $\{w_{j}\}_{j=1}^{4}$ on $\Omega=[0,1]\times(-1,1)$ for a $2\times 2$ decomposition.
@@ 2.1 FBPINN architecture
where $\pazocal{N}_{j}$ denotes a DNN associated with the $j$-th subdomain and parameterized by $p_{j}$ trainable parameters. To restrict the support of each local network $\pazocal{N}_{j}$ to its corresponding subdomain $\Omega_{j}$ and to appropriately weight the overlap between subdomains, we introduce a collection of “window” functions $\{w_{j}\}_{j=1}^{n_{s}}:\Omega\to\mathbb{R}$ such that $\operatorname{supp}(w_{j})\subset\Omega_{j}$ for all $j\in\{1,\dots,n_{s}\}$ and $\sum_{j=1}^{n_{s}}w_{j}\equiv 1$ on $\Omega$; this means that they form a partition of unity, see also Figure 1. Then, we define the global approximation space as $\pazocal{V}:=\sum_{j=1}^{n_{s}}w_{j}\,\pazocal{V}_{j}$. The solution $u$ is then approximated by a global network $\pazocal{N}$, obtained by combining the predictions of all subnetworks, i.e.,
@@
Here, the global parameter vector is given by $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{n_{s}})$. We use restriction operator ${\boldsymbol{R}}_{j}:\mathbb{R}^{p}\to\mathbb{R}^{p_{j}}$ to extract $\boldsymbol{\theta}_{j}$ from $\boldsymbol{\theta}$, i.e., $\boldsymbol{\theta}_{j}={\boldsymbol{R}}_{j}\boldsymbol{\theta}$. Conversely, ${\boldsymbol{R}}_{j}^{\top}$ is used to assign local parameters to the global network, i.e., $\boldsymbol{\theta}=\sum_{j}{\boldsymbol{R}}_{j}^{\top}\boldsymbol{\theta}_{j}$. Note, there is no overlap between the parameters of different subnetworks, see also Figure 2. Furthermore, to mitigate the spectral bias, the normalization function $\operatorname{norm}_{j}:\Omega\to(-1,1)^{d}$ is used to rescale the input coordinates on $\Omega_{j}$ in order to map high-frequency features in the global domain to lower-frequency representations on each subdomain. The FBPINN is then trained by inserting (6) into the loss function (4); cf. Moseley et al. (2023); Dolean et al. (2024).
@@ 3.1 Nonlinearly preconditioned LBFGS
$\boldsymbol{\theta}_{j}^{(\ast)}=\arg\min_{\boldsymbol{\theta}_{j}}\pazocal{L}_{j}(\boldsymbol{\theta}_{j};\pazocal{D}_{j}),$ (11)
with the local loss function
$\pazocal{L}_{j}(\boldsymbol{\theta}_{j};\pazocal{D}_{j})\coloneq\frac{1}{|\pazocal{D}_{j}|}\sum_{\vec{x}_{i}\in\pazocal{D}_{j}}\Big(\pazocal{P}\big[w_{j}(\vec{x})\,\pazocal{N}_{j}\big(\boldsymbol{\theta}_{j};\operatorname{norm}_{j}(\vec{x}_{j})\big)\big]-f(\vec{x}_{i})\Big)^{2},$
where $w_{j}$ corresponds to hard enforcement of homogeneous Dirichlet boundary conditions on $\partial\Omega_{j}$. The set of interior collocation points is denoted by $\pazocal{D}_{j}\subset\Omega_{j}$.
Equation 10 corresponds to a nonlinear additive Schwarz preconditioner in parameter space. In practice, the local problem (11) does not need to be solved exactly. Instead, an approximate solution suffices and can be obtained, for instance, by performing a fixed number $\eta$ of L-BFGS iterations initialized at $\boldsymbol{\theta}_{j}^{(k)}={\boldsymbol{R}}_{j}\boldsymbol{\theta}^{(k)}$. Here, we remark that local and global LBFGS maintain separate secant memories. The resulting right-preconditioned LBFGS algorithm is summarized in Algorithm 1.
@@
7:  $\boldsymbol{\theta}^{(k+1/2)}\leftarrow\boldsymbol{\theta}^{(k)}+\sum_{j=1}^{n_{s}}{\beta}^{(k)}_{j}\mathbf{R}_{j}^{\intercal}\mathbf{c}_{j}$ $\triangleright$ Compute $\boldsymbol{\beta}^{(k)}$ using approaches from Section 3.2
@@ 3.2 Scaling of Subdomain Corrections
We emphasize that the choice of the scaling parameters $\boldsymbol{\beta}$ is critical due to the intrinsic structure of FBPINNs and their differences from classical DD methods for discretized PDEs. Although the subnetworks do not share parameters (zero overlap), overlap in the spatial domain induces indirect coupling through their outputs, which causes the desynchronization of the solutions during the preconditioning step. In this work, we investigate three strategies for selecting the scaling parameters.
@@
  • Uniform scaling (UniS): As a first approach, we consider uniform scaling by a constant $\beta_{0}$, i.e., $\beta_{j}=\beta_{0}$, for all $j=1,\ldots,n_{s}$. The value of $\beta_{0}$ is typically set to $1$ in non-overlapping DD methods, but it can be fine-tuned. This approach has a low computational cost and is naturally parallel. However, it has been shown in Lee et al. that it does not yield an efficient preconditioner for NN problems.
  • Multiplicative line search scaling (LSS): Following Lee et al. , we also investigate a line search approach, where we start with a search direction $\mathbf{c}_{1}$ on the first subdomain, and look for the biggest value $\beta_{1}$ that satisfies $\pazocal{L}(\mathbf{\boldsymbol{\theta}}+\beta_{1}\mathbf{R}_{1}^{\intercal}\mathbf{c}_{1})<\pazocal{L}(\mathbf{\boldsymbol{\theta}})$. We then iterate through all search directions (subdomains) and seek their corresponding scaling factors such that
@@
Table 1 compares the estimated parallel computational costs of LBFGS and MP-LBFGS. Here, the word parallel refers to the per-device computational cost assuming a parallel implementation of FBPINN (one subdomain, one subnetwork per device), LBFGS, and MP-LBFGS. Notably, communication cost is not included in this estimate. For LBFGS, each epoch involves one forward and one backward pass to evaluate the gradient. In addition, the line search in (7) requires $\text{its}_{\text{ls}}$ forward evaluations. The computational cost of the LBFGS update scales with the number of stored secant pairs $Q$ as $2p+4Qp$, while the memory requirements accounting for the parameters, secant pairs, and momentum terms scale as $2p+2Qp$.
@@ 4.1 Convergence properties of multi-preconditioned LBFGS
First, we investigate the impact of the scaling strategies discussed in Section 3.2 on the performance of the MP-LBFGS method. Figure 3 shows that, independently of the chosen decomposition, the UniS strategy leads to unstable training. We further observe that as the number of subdomains increases (e.g., from $2\times 2$ to $4\times 4$), the SPM strategy significantly outperforms the LSS approach. This behavior can be attributed to the increasing distance between successive iterates after each line-search step. We also observe that, even for a fixed number of subdomains, the choice of decomposition strongly affects MP-LBFGS performance. For example, moving from a $2\times 2$ to a $4\times 1$ decomposition significantly degrades the LSS method, whereas the SPM strategy achieves an error reduction of more than one order of magnitude.
@@
Next, we compare the performance of MP-LBFGS with standard LBFGS. Results are reported in terms of epochs and effective parallel gradient evaluations $(\#g_{e})$. The metric $g_{e}$ approximately reflects the per-device computational work, while excluding communication overhead, costs of parameter updates and loss evaluations (see Table 1). The number of epochs serves as a proxy for communication cost.
We first examine the performance of MP-LBFGS while varying the number of local LBFGS iterations $\eta$, with $\eta=0$ corresponding to standard LBFGS. As shown in Figure 4, all MP-LBFGS configurations converge significantly faster in terms of epochs, thereby reducing communication cost. Moreover, MP-LBFGS requires a comparable, and in several cases lower, number of effective gradient evaluations $(\#g_{e})$. We also highlight, that for several values of $\eta$, MP-LBFGS achieves up to an order-of-magnitude reduction in the validation error $L_{2}^{\mathrm{val}}$.
Second, we compare LBFGS and MP-LBFGS with $\eta=5$ on the 2D Poisson and Burgers problems using $2\times 2$ and $3\times 3$ decompositions, while keeping the number of collocation points per subdomain fixed. As shown in Figure 5, MP-LBFGS consistently reduces the number of epochs while requiring comparable $g_{e}$ and achieving comparable accuracy across different decompositions. The only exception is the Burgers problem with a $3\times 3$ decomposition, where LBFGS attains higher accuracy; this outlier will be investigated in future work.
Figure 5: Convergence of MP-LBFGS (solid lines, $\eta=5$) and LBFGS (dotted lines) for Burgers (top) and Poisson (bottom) problem.
@@ 5 Conclusion
In this work, we developed a multi-preconditioned LBFGS (MP-LBFGS) framework for training FBPINNs. The proposed approach exploits the FBPINN architecture to compute local LBFGS search directions in parallel, while a global right-preconditioned LBFGS iteration ensures consistent convergence of the global minimization problem. A key component of the method is a novel subspace minimization strategy (SPM) for scaling locally computed search directions. Numerical experiments suggest that the proposed MP-LBFGS can achieve faster convergence, and higher accuracy than the standard LBFGS, while maintaining lower communication overhead. Future work will focus on parallel implementation, the incorporation of coarse spaces, and reducing the computational cost of solving the SPM subproblem.

### v2

@@
[…]
pdflscape
11institutetext: Marc Salvadó-Benasco 22institutetext: Università della Svizzera Italiana, 22email: marc.salvado@usi.ch 33institutetext: Aymane Kssim 44institutetext: Toulouse INP-ENSEEIHT, IRIT, ANITI; 44email: aymane.kssim@toulouse-inp.fr 55institutetext: Alexander Heinlein 66institutetext: Delft Institute of Applied Mathematics, TU Delft; 66email: a.heinlein@tudelft.nl 77institutetext: Rolf Krause 88institutetext: King Abdullah University of Science and Technology; 88email: rolf.krause@kaust.edu.sa 99institutetext: Serge Gratton 1010institutetext: Toulouse INP-ENSEEIHT, IRIT, ANITI; 1010email: serge.gratton@toulouse-inp.fr 1111institutetext: Alena Kopaničáková 1212institutetext: Toulouse INP-ENSEEIHT, IRIT, ANITI; 1212email: alena.kopanicakova@toulouse-inp.fr
@@ 1 Introduction
This manuscript is organized as follows. Sect. 2 introduces the FBPINN architecture, Sect. 3 presents the multi-preconditioned LBFGS algorithm, and numerical results and conclusions are given in Sect. 4 and Sect. 5, respectively.
@@ 2.1 FBPINN architecture
where $\pazocal{N}_{j}$ denotes a DNN associated with the $j$-th subdomain and parameterized by $p_{j}$ trainable parameters. To restrict the support of each local network $\pazocal{N}_{j}$ to its corresponding subdomain $\Omega_{j}$ and to appropriately weight the overlap between subdomains, we introduce a collection of “window” functions $\{w_{j}\}_{j=1}^{n_{s}}:\Omega\to\mathbb{R}$ such that $\operatorname{supp}(w_{j})\subset\Omega_{j}$ for all $j\in\{1,\dots,n_{s}\}$ and $\sum_{j=1}^{n_{s}}w_{j}\equiv 1$ on $\Omega$; this means that they form a partition of unity. Then, we define the global approximation space as $\pazocal{V}:=\sum_{j=1}^{n_{s}}w_{j}\,\pazocal{V}_{j}$. The solution $u$ is then approximated by a global network $\pazocal{N}$, obtained by combining the predictions of all subnetworks, i.e.,
@@
Here, the global parameter vector is given by $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{n_{s}})$. We use restriction operator ${\boldsymbol{R}}_{j}:\mathbb{R}^{p}\to\mathbb{R}^{p_{j}}$ to extract $\boldsymbol{\theta}_{j}$ from $\boldsymbol{\theta}$, i.e., $\boldsymbol{\theta}_{j}={\boldsymbol{R}}_{j}\boldsymbol{\theta}$. Conversely, ${\boldsymbol{R}}_{j}^{\top}$ is used to assign local parameters to the global network, i.e., $\boldsymbol{\theta}=\sum_{j}{\boldsymbol{R}}_{j}^{\top}\boldsymbol{\theta}_{j}$. Note, there is no overlap between the parameters of different subnetworks, see also Fig. 1. Furthermore, to mitigate the spectral bias, the normalization function $\operatorname{norm}_{j}:\Omega\to(-1,1)^{d}$ is used to rescale the input coordinates on $\Omega_{j}$ in order to map high-frequency features in the global domain to lower-frequency representations on each subdomain. The FBPINN is then trained by inserting (6) into the loss function (4); cf. Moseley et al. (2023); Dolean et al. (2024).
@@ 3.1 Nonlinearly preconditioned LBFGS
$\boldsymbol{\theta}_{j}^{(\ast)}=\arg\min_{\boldsymbol{\theta}_{j}}\pazocal{L}(\boldsymbol{\theta};\pazocal{D}),$ (11)
where $\boldsymbol{\theta}_{j}$ denotes the set of trainable parameters of the j-th subdomain.
Eq. (10) corresponds to a nonlinear additive Schwarz preconditioner in parameter space. In practice, the local problem (11) does not need to be solved exactly. Instead, an approximate solution suffices and can be obtained, for instance, by performing a fixed number $\eta$ of L-BFGS iterations initialized at $\boldsymbol{\theta}_{j}^{(k)}={\boldsymbol{R}}_{j}\boldsymbol{\theta}^{(k)}$. Here, we remark that local and global LBFGS maintain separate secant memories. The resulting right-preconditioned LBFGS algorithm is summarized in Alg. 1.
@@
7:  $\boldsymbol{\theta}^{(k+1/2)}\leftarrow\boldsymbol{\theta}^{(k)}+\sum_{j=1}^{n_{s}}{\beta}^{(k)}_{j}\mathbf{R}_{j}^{\intercal}\mathbf{c}_{j}$ $\triangleright$ Compute $\boldsymbol{\beta}^{(k)}$ using approaches from Sect. 3.2
@@ 3.2 Scaling of Subdomain Corrections
We emphasize that the choice of the scaling parameters $\boldsymbol{\beta}$ is critical due to the intrinsic structure of FBPINNs and their differences from classical DD methods. In particular, classical DD methods, originally proposed for elliptic PDEs, exploit structural properties of finite-element/difference discretizations, such as locality of basis functions and sparse, spatially localized coupling between degrees of freedom. This ensures that subdomain corrections mainly affect local error components, while global modes are handled through coarse-grid or global synchronization steps.
In contrast, machine-learning problems involve nonconvex minimization, for which only a few domain-decomposition methods have been developed, e.g., Gratton et al. (2025). Moreover, locally defined physical degrees of freedom are replaced by globally coupled parameters, fundamentally altering error-propagation mechanisms. As a result, classical DD assumptions, such as convexity, locality, and spectral separation, no longer hold, which often leads to degraded convergence of conventional DD methods.
For the FBPINNs considered in this work, spatial overlap induces indirect coupling, leading to desynchronization of parameters during the preconditioning step and motivating the design of novel scaling strategies for the subdomain corrections. In particular, we investigate three such strategies:
@@
  • Uniform scaling (UniS): As a first approach, we consider uniform scaling by a constant $\beta_{0}$, i.e., $\beta_{j}=\beta_{0}$, for all $j=1,\ldots,n_{s}$. The value of $\beta_{0}$ is typically set to $1$ in non-overlapping DD methods, but it can be fine-tuned. This approach has a low computational cost and is naturally parallel. However, it has been shown in Lee et al. (2026) that it does not yield an efficient preconditioner for NN problems.
  • Multiplicative line search scaling (LSS): Following Lee et al. (2026), we also investigate a line search approach, where we start with a search direction $\mathbf{c}_{1}$ on the first subdomain, and look for the biggest value $\beta_{1}$ that satisfies $\pazocal{L}(\mathbf{\boldsymbol{\theta}}+\beta_{1}\mathbf{R}_{1}^{\intercal}\mathbf{c}_{1})<\pazocal{L}(\mathbf{\boldsymbol{\theta}})$. We then iterate through all search directions (subdomains) and seek their corresponding scaling factors such that
@@
Tab. 1 compares the estimated parallel computational costs of LBFGS and MP-LBFGS. Here, the word parallel refers to the per-device computational cost assuming a parallel implementation of FBPINN (one subdomain, one subnetwork per device), LBFGS, and MP-LBFGS. Notably, communication cost is not included in this estimate. For LBFGS, each epoch involves one forward and one backward pass to evaluate the gradient. In addition, the line search in (7) requires $\text{its}_{\text{ls}}$ forward evaluations. The computational cost of the LBFGS update scales with the number of stored secant pairs $Q$ as $2p+4Qp$, while the memory requirements accounting for the parameters, secant pairs, and momentum terms scale as $2p+2Qp$.
@@ 4.1 Convergence properties of multi-preconditioned LBFGS
First, we investigate the impact of the scaling strategies discussed in Sect. 3.2 on the performance of the MP-LBFGS method. Fig. 2 shows that, independently of the chosen decomposition, the UniS strategy leads to unstable training. We further observe that as the number of subdomains increases (e.g., from $2\times 2$ to $4\times 4$), the SPM strategy significantly outperforms the LSS approach. This behavior can be attributed to the increasing distance between successive iterates after each line-search step. We also observe that, even for a fixed number of subdomains, the choice of decomposition strongly affects MP-LBFGS performance. For example, moving from a $2\times 2$ to a $4\times 1$ decomposition significantly degrades the LSS method, whereas the SPM strategy achieves an error reduction of more than one order of magnitude.
@@
Next, we compare the performance of MP-LBFGS with standard LBFGS. Results are reported in terms of epochs and effective parallel gradient evaluations $(\#g_{e})$. The metric $g_{e}$ approximately reflects the per-device computational work, while excluding communication overhead, costs of parameter updates and loss evaluations (see Tab. 1). The number of epochs serves as a proxy for communication cost. Here, we also note that, for all presented experiments, the number of Newton iterations in the SPM method, denoted as $\text{its}_{\text{Newton}}$, is on average equal to two.
We first examine the performance of MP-LBFGS while varying the number of local LBFGS iterations $\eta$, with $\eta=0$ corresponding to standard LBFGS. As shown in Fig. 3, all MP-LBFGS configurations converge significantly faster in terms of epochs, thereby reducing communication cost. Moreover, MP-LBFGS requires a comparable, and in several cases lower, number of effective gradient evaluations $(\#g_{e})$. We also highlight, that for several values of $\eta$, MP-LBFGS achieves up to an order-of-magnitude reduction in the validation error $L_{2}^{\mathrm{val}}$.
Second, we compare LBFGS and MP-LBFGS with $\eta=5$ on the 2D Poisson and Burgers problems using $2\times 2$ and $3\times 3$ decompositions, while keeping the number of collocation points per subdomain fixed. As shown in Fig. 4, MP-LBFGS consistently reduces the number of epochs while requiring comparable $g_{e}$ and achieving comparable accuracy across different decompositions. Note that, here, one epoch of the MP-LBFGS algorithm with $\eta=5$ corresponds to seven gradient evaluations; see Tab. 1 for details. The only exception is the Burgers problem with a $3\times 3$ decomposition, where LBFGS attains higher accuracy. Our empirical evidence suggests that, for this particular example, the non-convexity of the problem causes MP-LBFGS to converge to a local minimum with a larger error than L-BFGS.
Figure 4: Convergence of MP-LBFGS (solid lines, $\eta=5$) and LBFGS (dotted lines) for Burgers (top) and Poisson (bottom) problem. Experiments terminate after $g_{e}$ of $20,000$ or $40,000$ is reached.
@@ 5 Conclusion
In this work, we developed a multi-preconditioned LBFGS (MP-LBFGS) framework for training FBPINNs. The proposed approach exploits the FBPINN architecture to compute local LBFGS search directions in parallel, while a global right-preconditioned LBFGS iteration ensures consistent convergence of the global minimization problem. A key component of the method is a novel subspace minimization strategy (SPM) for scaling locally computed search directions. Numerical experiments suggest that the proposed MP-LBFGS can achieve faster convergence, and higher accuracy than the standard LBFGS, while maintaining lower communication overhead. Future work will focus on parallel implementation, allowing for wall-clock comparison and the incorporation of more complex examples as well as on incorporation of coarse spaces; and on reducing the cost of solving the SPM subproblem.